Leah, Inc., is proposing a rights offering. Presently there are 200,000 shares outstanding at $45 each. There will be 25,000 new shares offered at $40 each.

a. What is the new market value of the company? (Do not round intermediate calculations.)
b. How many rights are associated with one of the new shares? (Do not round intermediate calculations.)
c. What is the ex-rights price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
d. What is the value of a right? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

a)

New market value = Existing shares x price per share + rights offering x price per share

New market value =(200000*45)+(25000*40)

New market value =10000000

b)

Number of shares needed for one right share = outstanding shares/rights share issued

Number of shares needed for one right share =200000/25000

Number of shares needed for one right share =8

c)

Ex-rights per share = new market value of the company/(existing shares + rights offerings)

Ex-rights per share =10000000/(200000+25000)

Ex-rights per share =44.44

d)

Value of the right = The price of the outstanding share – the new price of the share

Value of the right =45-44.44

Value of the right =0.56